Microbiology

Microbiology is the study of tiny micro-organisms. For a long time, it was not known that they existed, because microscopes had not yet been invented. Micro-organisms can be one-celled or multi-celled organisms. They include fungi (plural of fungus), bacteria (plural of bacterium), and viruses. (Viruses are not strictly classified as living organisms because of the technical definitions about what living organisms are).

Hence, microbiology began after the invention of the microscope, especially what is called the compound microscope. The name is given because you can see organisms of the size of 1 micrometer (or 1 micron) with such a device. One micron is 1/10,00,000 meter or 0.001 millimeter. For comparison, thick black human hair is about 0.1 millimeter or 100 microns. Most plant and animal cells and their composition can be seen with such microscopes. All creatures studied in micro-biology are called microbes.

¡compound microscope picture¿

Microbes

Microbes are among the most diverse forms occurring in nature. Since they are so small and hard to see, only a few microbes have been well-studied or even isolated (discovered) in the laboratory. Micro-biologists, who study microbes, believe that only about 1microbes have been studied!

Mostly, those microbes that influence or affect humans or their lives and life-stock (cattle, pets, crops, etc) have been studied in detail. As sophisticated machines get built, more and more precise information is becoming available about these microbes. Hence the field of microbiology is a growing and important one.

It is important to remember that microbes can be both very helpful or harmful to humans.

Fungus

Yeast, which causes bread to rise and become fluffy and soft, is a type of fungus. Moulds and mushrooms are also types of fungi. Some mushrooms are poisonous, so it is important to know which ones can be eaten.

Robert Hooke and Antonie van Leeuwenhoek were among the first to observe micro-organisms using a microscope. Robert Hooke was the first to study moulds, which form the growth you see on bread, for example.

¡bread figure¿ Bacteria Not all bacteria cause disease, although bacteria were first discovered in relation to diseases. Louis

Pasteur realised in the 19th century that ”germs” cause disease and proposed the method now called pateurisation to kill mould and bacteria in milk. He also studied the bacteria that caused cholera in chickens.

Pasteur learned that when a weakened form of the bacteria was given to healthy people, they got mild forms of usually deadly diseases. This led to the idea of vaccination, which is very common for several diseases today. In fact, the bacteria that cause smallpox is now completely eradicated (no longer exists in sufficient strength in the natural state) and children are no longer even vaccinated against this once-deadly disease.

Later, Robert Koch discovered the tuberculosis bacillus. He also highlighted the importance of sterilising instruments in surgery and medicine so that new patients were not infected with the disease of earlier patients on whom the equipment had been used.

Typhoid disease is a water-borne disease that causes diarrhoea, fever, and even death. But anti-biotics have proven very effective to treat this. Did you know that typhoid-causing bacteria (called Salmonella typhi) can only live in humans. So, for the bacteria to survive, they have to infect as many people as possible! By maintaining hygiene and being careful of what you eat and drink, typhoid can be prevented in your entire neighbourhood.

¡typhi¿

In modern times, food poisoning through contamination by bacteria called E. coli (Escherichia coli) is very common. However, not all E. coli are dangerous. Several of them live in human intestines and help in digestion. While some of them help to process waste matter, others help in absorbing nutrients from the digested food. E. coli are commonly found in microbiology labs for experiments since they already exist in humans and so are harmless. A dish of E. coli double in number in about half an hour, so they are very convenient to work with.

¡E coli figure¿

Viruses

Viruses are much smaller than bacteria. While many of them may not directly cause disease, many of the most dangerous diseases in the world today are caused by viruses. Also, anti-biotics are not effective against them. Most viral infections can only be controlled and not treated.

For instance, avian influenza (bird flu) and foot-and-mouth disease in cattle which are caused by viruses recently affected large parts of the world. In both cases, the spread of the disease was halted by killing all the animals who were suspected to have been infected. SARS (Severe acute respiratory syndrome) that killed many hundreds of humans in 2002-2003 also is caused by viruses. No known treatment exists today, including for the HIV virus.

Milder diseases caused by viruses are chicken pox and the common cold. It is said that ”it takes 7 days to recover from cold without medication and 1 week to recover with medication” which may well be true!

Viruses have different mechanisms by which they cause disease. This is the main difficulty with treating diseases caused by viruses. Also, viruses themselves keep on (genetically) changing; this is called mutation and makes them hard to cure (a drug that cures the original variety may not be effective against the mutant).

Understanding human disease from viruses is a subject of central importance in today’s world.

BOX

Robert Hooke’s book Micrographia describes some mould found on a book cover.

”But, first, I must premise a short description of this Specimen, which is nothing else but the appearance of a small white spot of hairy mould, multitudes of which I found to bespeck whiten over the red covers of a small book, which, it seems, were of Sheeps skin, that being more apt to gather mould, even in a dry and clean room, then other leathers. These spots appear’d, through a good Microscope, to be a very pretty shap’d Vegetative body, which, from almost the same part of the Leather, shot out multitudes of small long cylindrical and transparent stalks, not exactly streight, but a little bended with the weight of a round and white knob that grew on the top of each of them; many of these knobs I observ’d to be very round, and of a smooth surface, others smooth likewise, but a little oblong; several of them a little broken, or cloven with chops at the top; others flitter’d as ’twere, or flown all to pieces. The whole substance of these pretty bodies was of a very tender constitution, much like the substance of the softer kind of common white Mushroms.”